### v1

# S3-Attention:Attention-Aligned Endogenous Retrieval for Memory-Bounded Long-Context Inference

Abstract
Large language models are increasingly used on multi-document and long-form inputs, yet long-context inference remains memory- and noise-inefficient: KV caching scales linearly with context length, and external retrievers may return lexically similar but causally irrelevant passages. We present S3-Attention, a memory-first inference-time framework that treats long-context usage as attention-aligned endogenous retrieval. S3-Attention decodes transient key/query projections into Top-k sparse feature IDs via lightweight sparse autoencoders and builds a CPU inverted index (feature → token positions/spans) during a streaming scan, allowing the KV cache to be discarded and keeping GPU memory bounded by the scan chunk size. At generation time, feature co-activation retrieves compact evidence spans, optionally fused with BM25 for exact matching. Under a unified LongBench evaluation protocol (fixed prompting/decoding and matched token budgets), S3-Hybrid closely matches full-context inference on multiple model families and improves robustness on several information-dense settings. We also report the current engineering limitation that our prototype incurs higher wall-clock latency than optimized FullKV baselines, motivating future kernel-level optimization.
Machine Learning, ICML
1 Introduction
Large language models (LLMs) have become a default interface for question answering, summarization, and multi-step reasoning. As these systems move from single documents to collections of reports, codebases, and conversational histories, they increasingly need to condition on contexts that far exceed a single attention window(Ding et al., 2024; Zhao et al., 2024a). The prevailing response is to scale context lengths through continued training or architectural modifications(Liu et al., 2025b; Chen et al., 2024b; Hu et al., 2024; Du et al., 2025; Mao et al., 2025; Mohtashami and Jaggi, 2023), with recent models advertising windows from 128K up to million-token regimes.
However, longer context windows do not automatically translate into reliable long-context reasoning. Full-context inference remains deployment-unfriendly(Ma et al., 2024): self-attention incurs quadratic compute(Lou et al., 2024), and KV caches scale(Sun et al., 2024) linearly with the number of processed tokens(Tang et al., 2024), quickly saturating GPU memory(Wang et al., 2025a). At the same time, real-world long inputs are typically sparse in signal—only(Zhu et al., 2025) a small fraction of tokens are causally useful for answering a given query—so naively attending to everything can dilute evidence and amplify distraction(Xu et al., 2024; Hooper et al., 2025; Zarch et al., 2025).
A pragmatic alternative is Retrieval-Augmented Generation (RAG)(Lewis et al., 2020; Cheng et al., 2025), which replaces full-context attention with an external retriever that selects a small set of passages for the generator. While RAG is memory-efficient and widely used, it introduces a semantic mismatch: retrieval is performed in an embedding space whose similarity structure need not align with the generator’s internal reasoning features(Wei et al., 2025). In noisy or heterogeneous corpora, this mismatch can surface lexically overlapping yet causally irrelevant text, degrading multi-hop reasoning(Liu et al., 2025a) and increasing hallucinations(Wang et al., 2025b).
These limitations motivate a complementary direction: endogenous retrieval, where the model retrieves evidence using its own internal signals(Wu et al., 2024). Recent results support this premise(Zhang et al., 2025). For instance, InfiniRetri links attention allocation patterns to retrieval behavior and leverages attention information to retrieve needles from extremely long inputs without additional training (Ye et al., 2025). Meanwhile, KV-cache compression and sparse-attention methods use attention scores(Xiao et al., 2024a) or other internal statistics to decide which past states to retain. Yet directly operating on dense attention weights or cached states does not provide a scalable, queryable memory(Chen et al., 2025b): the representations are continuous, high-dimensional, and expensive to store or search at token granularity(Ma et al., 2024; Liu et al., 2024, 2025c), and most systems are optimized for efficiency rather than explicitly surfacing interpretable evidence spans.
We propose S3-Attention (Sparse & Semantic Streaming Attention), which turns long-context inference into a streaming, attention-aligned retrieval procedure. The key idea is to compress the transient key and query projections—already computed by the model—into discrete sparse semantic features that can be indexed efficiently. Concretely, we train lightweight Top-$k$ sparse autoencoders (SAEs) on attention projections and use them at inference time to map each token span to a small set of active features. During a single streaming prefill pass, S3-Attention encodes and stores an inverted index (feature $\rightarrow$ spans) on CPU and then discards the KV cache, achieving O(1) GPU memory with respect to total context length. At generation time, the query’s SAE features retrieve high-density spans via feature co-activation, which we refine with density smoothing and non-maximum suppression. To improve robustness on exact entity matching and other lexical cues, we optionally fuse the endogenous score with BM25, yielding S3-Hybrid.
We evaluate S3-Attention on LongBench across multiple LLM families and tasks in a unified inference environment. S3-Hybrid retains 99.4% of full-context performance on Llama-3-8B (24.87 vs. 25.01) and over 99% on Qwen2-7B (18.80 vs. 18.92), while enabling constant-GPU-memory long-context processing. Beyond average scores, we observe consistent gains on multi-document and multi-hop settings and a “denoising” effect on information-dense tasks where selectively filtering context can outperform full-context attention.
Contributions.
This paper makes three contributions:
  • We articulate long-context inference as an endogenous retrieval problem and motivate why aligning retrieval with internal attention representations can mitigate the semantic gap of exogenous retrievers.
  • We introduce S3-Attention, which uses SAE-decoded sparse semantic features to build a streaming inverted index and achieve O(1) GPU memory without fine-tuning the base LLM.
  • We demonstrate strong empirical performance on LongBench and show that fusing endogenous semantic signals with BM25 yields a robust hybrid retriever with near-lossless fidelity to full-context inference.
S3-Attention is designed for settings where GPU memory is the primary bottleneck and only a small fraction of a long input is causally useful for a given query. Our goal is not to outperform all KV-compression or advanced RAG pipelines in wall-clock latency, but to provide a queryable, attention-aligned indexing mechanism that (i) avoids storing dense KV states, (ii) retrieves interpretable evidence spans, and (iii) can be combined with lexical retrieval for robustness.
2 Related Works
Long-context reasoning with large language models (LLMs) involves a fundamental trade-off between fidelity, efficiency, and robustness. Feeding the full context avoids information loss but incurs prohibitive computation and KV-cache overhead, while irrelevant or noisy tokens increasingly dilute causally relevant evidence as context length grows. Prior work shows that only a small subset of past tokens materially contributes to the current prediction, motivating query-aware selection or sparsification in long-context settings (Hu et al., 2024; Tang et al., 2024). Existing methods can be broadly categorized into approaches based on exogenous versus endogenous signals.
Exogenous signals: Retrieval-Augmented Generation (RAG) is the canonical exogenous approach, where an external retriever selects passages deemed relevant to the query (Lewis et al., 2020). Subsequent work explores system-level improvements such as reranking, query decomposition, and speculative decoding (Li et al., 2025; Liu et al., 2025d; Yang et al., 2025). However, RAG is highly sensitive to retrieval quality: in long or noisy contexts it may miss critical evidence or select spurious passages (Xian et al., 2024; Chen et al., 2024a). More fundamentally, relevance defined by external similarity does not necessarily align with the information the LLM internally uses for generation, leading to implicit misalignment (Jin et al., 2025). Reflection-based variants partially alleviate retrieval errors (Asai et al., 2024; Chen et al., 2025a), but remain dependent on external control signals.
Endogenous signals: Another line of work studies internal mechanisms of LLMs. In-context learning has been formalized as conditioned associative memory retrieval (Smart et al., 2025), and empirical analyses show that long-context localization often relies on a small number of retrieval-oriented attention heads (Zhao et al., 2024b). Query-aware sparsification and causal retrieval further indicate that only a limited subset of tokens affects prediction outcomes (Hu et al., 2024; Tang et al., 2024). Endogenous signals have also been used for KV-cache selection or compression to improve inference efficiency (Zhang et al., 2023; Li et al., 2024; Wu et al., 2025), but these methods primarily target system efficiency and do not explicitly identify interpretable, causally relevant evidence segments for answering long-context questions (Huang et al., 2025).
Figure 1: Overview of the $S^{3}$-Attention framework. The framework consists of two phases connected by a Top-$k$ Sparse Autoencoder (SAE). Streaming Semantic Indexing (red flow) encodes transient key projections into sparse semantic features to build a CPU-based inverted index, enabling the dense KV cache to be discarded and maintaining an $O(1)$ GPU memory footprint. Endogenous Retrieval (blue flow) encodes query projections with the same SAE, where activated features retrieve relevant context spans via feature co-activation, which are then fed back into the LLM for answer generation.SAE is trained on K projections and reused to discretize Q.
3 Methodology
3.1 Overview: From Exogenous to Endogenous Retrieval
Handling long contexts in Large Language Models (LLMs) presents a trilemma among fidelity (Full-Attention), efficiency (RAG), and robustness (Noise Tolerance). We argue that the noise sensitivity observed in standard Retrieval-Augmented Generation (RAG) stems from a fundamental Semantic Gap: the reliance on exogenous retrievers (e.g., BERT-based embeddings) whose latent spaces are misaligned with the reasoning heads of the generative LLM. This heterogeneity often leads to the retrieval of superficially similar but causally irrelevant segments, injecting noise that hallucinates the generation.
Drawing inspiration from human cognitive processes—specifically how readers engage with information-dense texts (e.g., news reports or technical manuals)—we observe that humans do not process every word with equal weight. Instead, they employ a goal-oriented selective attention mechanism, scanning the text to lock onto salient information relevant to their query while actively filtering out background noise.
To replicate this endogenous process in LLMs, we propose S3-Attention (Sparse & Semantic Streaming Attention). Bridging the gap between memory constraints and context understanding, S3-Attention replaces the external retriever with an Endogenous Retrieval mechanism. By decoding the model’s own transient Key and Query states into sparse semantic features, we enable the model to perform “introspection”—explicitly retaining only the context segments that trigger its own attention mechanisms. This approach ensures that the retrieved context is cognitively aligned with the generation process.
3.2 Theoretical Framework: The Endogenous Information Bottleneck
We view long-context compression as selecting a compact subset of the input that preserves task-relevant evidence for answering a query. Under this perspective, an ideal objective can be formulated using mutual information, for example by maximizing
$I(Y;\,\hat{C}\mid Q)$
subject to a compression constraint on $\hat{C}$. However, this quantity is intractable to estimate for modern large language models, and we do not attempt to compute it in practice.
Instead, we propose an operational endogenous proxy derived from the model’s own attention-matching signals. Specifically, we discretize transient attention projections into a sparse feature space and rank context positions by their inverse-document-frequency (IDF) weighted feature co-activation with the query. Appendix B provides a motivating inequality under explicit simplifying assumptions. We emphasize that this result should be interpreted as a heuristic justification of our scoring rule, rather than as a guarantee that feature overlap faithfully tracks mutual information in real models.
The Exogenous Gap. Standard RAG methods optimize a surrogate objective $\max\text{Sim}(\text{Embed}(c),\text{Embed}(Q))$ using external encoders (e.g., BGE, Contriever). However the external embedding space does not capture the LLM’s internal notion of “relevance to YY”.
Our Insight. We observe that the attention mechanism implicitly solves a related problem. Let $\mathbf{A}=\text{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d})$ be the attention matrix. High attention weights $A_{ij}$ indicate that token $j$ in the context is causally useful for predicting the next token at position $i$. By extracting the LLM’s own Key-Query matching patterns, we obtain an endogenous relevance signal that is inherently aligned with the generation process.
From Attention to Sparse Features. Direct use of attention weights is infeasible due to quadratic complexity. Instead, we leverage Sparse Autoencoders (SAEs) to decompose the dense Key/Query vectors into interpretable sparse features. We show in Appendix D that under strong simplifying assumptions, feature co-occurrence between $\mathbf{K}$ and $\mathbf{Q}$ provides a tractable lower bound on mutual information:
$I(Y;\hat{C}\mid Q)\geq\mathbb{E}\left[\sum_{t\in\hat{C}}\sum_{f\in\mathcal{F}_{Q}}\mathbb{1}[f\in\mathcal{F}_{t}]\cdot w_{f}\right]+\text{const}$ (1)
where $\mathcal{F}_{Q}$ and $\mathcal{F}_{t}$ are the active SAE features for the query and context token $t$, respectively. This directly motivates our scoring function in Eq. 7.
3.3 Deciphering Attention via Sparse Autoencoders
The foundation of our approach is the ability to interpret the polysemantic activation vectors within the LLM’s attention heads. We employ Top-K Sparse Autoencoders (SAEs) (Gao et al., 2024) to decompose these dense vectors into interpretable sparse features.
Let $\mathbf{x}\in\mathbb{R}^{d_{head}}$ be the activation vector projected by the Key ($\mathbf{K}$) or Query ($\mathbf{Q}$) matrix in layer $\ell$. An SAE consists of an encoder and a decoder. The encoder projects $\mathbf{x}$ into a higher-dimensional latent space $\mathbb{R}^{d_{latent}}$ (where $d_{latent}\gg d_{head}$):
$\mathbf{z}=\text{ReLU}(\mathbf{W}_{enc}(\mathbf{x}-\mathbf{b}_{dec})+\mathbf{b}_{enc})$ (2)
To enforce sparsity, we apply a Top-K nonlinearity, retaining only the $k$ most active latent features:
$\hat{\mathbf{z}}=\text{TopK}(\mathbf{z},k),\quad\|\hat{\mathbf{z}}\|_{0}=k$ (3)
The original activation is reconstructed as $\hat{\mathbf{x}}=\hat{\mathbf{z}}\mathbf{W}_{dec}+\mathbf{b}_{dec}$. We train a SAE on key projections and reuse the same SAE to encode query projections, ensuring a feature-ID space.
3.4 Phase 1: Streaming Semantic Indexing
To process Massive contexts without maintaining a linear-growth KV cache, we introduce a streaming indexing protocol.
We process the input context $\mathcal{C}$ in sequential chunks $\{c_{1},\dots,c_{m}\}$. For each chunk, the LLM performs a forward pass to generate Key states. Instead of caching these tensors, we immediately encode them via the SAEs into sparse indices:
$\mathcal{F}_{t}^{(\ell)}=\text{Indices}(\text{SAE}_{\ell}(\mathbf{k}_{t}^{(\ell)}))$ (4)
where $\mathcal{F}_{t}^{(\ell)}$ represents the set of active semantic features for the token at position $t$ in layer $\ell$.
We construct a lightweight Inverted Semantic Index $\mathcal{I}$ on the CPU, mapping each feature ID to a list of absolute positions:
$\forall f\in\mathcal{F}_{t}^{(\ell)},\quad\mathcal{I}_{\ell}[f]\leftarrow\mathcal{I}_{\ell}[f]\cup\{t\}$ (5)
Crucially, once the features are indexed, the GPU memory for activations and KV cache is released. This transforms the memory complexity of the pre-filling phase from $\mathcal{O}(L)$ to $\mathcal{O}(1)$, limited only by chunk size.
CPU Index Size.
S3-Attention achieves $O(1)$ GPU memory with respect to the total context length by discarding the KV cache after feature decoding, but it still maintains a CPU-side inverted index. Let $L$ denote the number of context tokens, $|L_{\text{target}}|$ the number of instrumented layers, and $k$ the Top-$k$ sparsity per token after head aggregation. The total number of postings is
$P=L\cdot|L_{\text{target}}|\cdot k.$
In an idealized implementation that stores token positions as int32, this corresponds to approximately $4P$ bytes for positions alone (e.g., $L=128\text{K},\ |L_{\text{target}}|=4,\ k=128\Rightarrow$ about $256\,\text{MiB}$). Our current prototype uses Python dict/list posting lists, which incur substantial overhead; production implementations should instead use contiguous integer arrays with delta encoding and optional stop-feature pruning to reduce constant factors.
3.5 Phase 2: Semantic Density Estimation & Retrieval
Upon receiving a query $Q$, our goal is to identify regions in the context that maximizes the expected attention of the LLM.
Query Decoding. We compute the Query projections for $Q$ and encode them using the same SAEs used for indexing. This extracts the model’s intrinsic “search intent”:
$\{(w_{q},f_{q})\}=\text{SAE}_{\ell}(\mathbf{q}^{(\ell)})$ (6)
where $w_{q}$ is the activation strength of feature $f_{q}$.
Semantic Density Estimation. We calculate a semantic relevance score $S[t]$ for every position $t$ in the context. Unlike vector similarity search which treats text as static blocks, we formulate this as a feature voting process:
$S[t]=\sum_{\ell\in\mathcal{L}_{target}}\sum_{f\in f_{q}^{(\ell)}}\mathbb{1}(t\in\mathcal{I}_{\ell}[f])\cdot w_{q,f}^{(\ell)}\cdot\text{IDF}(f)$ (7)
Here, $\text{IDF}(f)$ down-weights high-frequency features (e.g., common syntactic patterns) to focus on rare, information-rich concepts.
Adaptive Granularity. Fixed-size chunking (used in RAG) often fragments semantic units. S3-Attention employs a dynamic approach. We apply a 1D convolution kernel to smooth the sparse score array $S$, generating a semantic density curve. We then perform Non-Maximum Suppression (NMS) to identify peak density regions. For each peak, we retrieve a variable-length span, ensuring the context is cut at natural semantic boundaries rather than arbitrary token counts.
3.6 Phase 3: Hybrid Fusion
To guarantee robustness across diverse tasks, we fuse the endogenous semantic signal with structural and lexical priors:
$\mathcal{M}_{final}=\mathcal{M}_{S^{3}}\cup\mathcal{M}_{BM25}\cup\mathcal{M}_{Bias}$ (8)
  • $\mathcal{M}_{S^{3}}$: Indices selected by the SAE-driven endogenous retrieval, capturing abstract reasoning chains.
  • $\mathcal{M}_{BM25}$: Indices selected by lexical matching, compensating for rare entities (e.g., random IDs) that SAEs might not reconstruct perfectly.
  • $\mathcal{M}_{Bias}$: A positional bias retaining the first and last $N$ tokens (Lead/Tail), mitigating the “Lost-in-the-Middle” phenomenon.
The fused indices are used to gather the original tokens into a compressed context $\tilde{\mathcal{C}}$, which is fed to the LLM for generation.
4 Experiments
4.1 Experimental Setup and Methodology
To demonstrate the universality of S3-Attention, we conduct experiments across three distinct state-of-the-art open-weights LLMs (Llama-3.1, Mistral, and Qwen2) using bfloat16 precision. These models were selected to represent diverse attention mechanisms and positional embeddings. We evaluate performance using the LongBench suite, focusing on nine datasets spanning Single-Document QA, Multi-Document QA, and Summarization tasks. This selection rigorously tests the models’ ability to retrieve and synthesize information from contexts exceeding 10k tokens.
Our pipeline consists of two stages: Offline SAE Training and Online Inference. For the offline stage, we train Top-K Sparse Autoencoders (SAEs) on general linguistic corpora to create a feature space for efficient discretization. For online inference, we benchmark S3-Attention against strong baselines, including: (1) Full-Context (upper bound), (2) Retrieval-Augmented Generation (RAG), and (3) KV cache compression methods such as H2O (Zhang et al., 2023) and StreamingLLM (Xiao et al., 2024b). To ensure rigorous comparison, we employ a unified evaluation protocol with identical decoding settings and prompts across all methods.
Specific details regarding model versions, dataset breakdowns, SAE training configurations, and decoding parameters are provided in Appendix A.
4.2 Qualitative Analysis: Mechanism of Endogenous Alignment
To validate our hypothesis that endogenous retrieval eliminates the semantic gap inherent in exogenous methods, we conducted a microscopic analysis of the retrieval behaviors. Figure 2 visualizes the retrieval scores of standard RAG (BGE-Small) against the semantic activation maps of S3-Attention on a multi-hop reasoning query: “Which film starring Tom Hanks was directed by Steven Spielberg?”
Visualizing the Semantic Gap.
As shown in Figure 2, the contrast between the two paradigms explains the performance gap observed in our main experiments:
  • Exogenous RAG Failure (Surface-Level Matching): The RAG retriever (Top Panel) falls into a “lexical trap.” It assigns the highest similarity score ($0.751$) to Sentence 1 (green bar), which is a generic biography of Tom Hanks. While this sentence shares high lexical overlap with the query entities (“Tom Hanks”), it contributes zero informational value to the specific question. The actual answer (Sentence 5, “The Post”) is ranked lower ($0.635$), burying the true signal under noise. This confirms that embedding models often prioritize topical similarity over truthfulness.
  • Endogenous S3 Success (Deep Semantic Anchoring): In stark contrast, S3-Attention (Bottom Panel) demonstrates an intrinsic ability to filter out noise. The SAE-decoded semantic activations are near-zero for the generic biography section. Instead, we observe sharp, high-confidence activation peaks (marked by red arrows) precisely at the tokens “The Post” and conceptually related terms like “Pentagon Papers.” This indicates that the LLM is not merely matching names; it is attending to the causal evidence required to resolve the query. The mechanism effectively acts as a semantic band-pass filter, suppressing the “Tom Hanks” biography noise while amplifying the specific film entity.
This phenomenon is consistent across our dataset. S3-Attention achieves a superior Signal-to-Noise Ratio (SNR), activating only the 1–2% of tokens that serve as reasoning bridges, whereas RAG retrieves broad, lexically dense but logically shallow chunks.
Figure 2: Endogenous vs. Exogenous Retrieval. Top: RAG (BGE-Small) is distracted by high lexical overlap, ranking a generic biography (Sentence 1) higher than the true answer (Sentence 5). Bottom: S3-Attention (Ours) ignores the distraction, showing sparse activation peaks exclusively at the semantic answer anchor (“The Post”) and its reasoning evidence (“Pentagon Papers”).
Additional visualization results and analysis of other datasets are provided in Appendix F.
4.3 Quantitative Results on LongBench
Performance retention is defined as Score(method) / Score(FullKV) under the same prompt template, decoding parameters, and evaluation script. All “near-lossless” statements in this paper refer to retention under this unified protocol.
We present the comprehensive evaluation results on LongBench in Table 3. It is important to note that absolute evaluation scores on LongBench can fluctuate significantly depending on inference environments (e.g., prompt templates and quantization kernels). To ensure a rigorous comparison, we distinguish between standard baselines (provided for reference) and methods evaluated within our unified environment (marked in bold), which include FullKV, RAG, BM25, and our S3-Attention variants.
Performance Retention: Near-Lossless Compression.
Instead of merely comparing absolute metrics across disparate environments, we focus on the Performance Retention Rate relative to the FullKV upper bound.
  • Superior Fidelity: While reference baselines like SnapKV achieve high absolute scores in lenient environments, they typically exhibit a performance drop of 7–10% compared to their corresponding full-context baselines (e.g., SnapKV 30.61 vs. Ref-FullKV 32.92 on Llama-3). In contrast, within our strictly unified setting, S3-Hybrid demonstrates exceptional fidelity. For Llama-3-8B, it achieves an average score of 24.87, retaining 99.4% of the internal FullKV performance (25.01).
  • Stability Across Models: This trend holds for Qwen2-7B, where S3-Hybrid (18.80) matches the FullKV baseline (18.92) within the margin of error (>>99% retention), significantly outperforming the sparse retrieval baseline BM25 (18.80 vs 18.80, note: tied in avg but S3 is higher in Multi-Doc). …demonstrating that our SAE-driven indexing introduces negligible information loss compared to the theoretical upper bound.
The ”Denoising” Effect in Information-Dense Tasks.
Although FullKV generally serves as an upper bound, S3-Attention exhibits a counter-intuitive ”Less is More” phenomenon on specific tasks requiring precise evidence extraction. On Qasper (Llama-3), a paper reading task, S3-Hybrid scores 21.50, surpassing both the FullKV baseline (20.56) and the exogenous RAG baseline (21.43). We attribute this to the Semantic Band-Pass Filter effect: by actively pruning irrelevant context via SAE feature selection, our method reduces the distraction noise that often confuses the model during full-context attention, effectively acting as a cleaner signal source than the raw document.
4.4 Comparison with Exogenous Retrieval (RAG)
A central question is how Endogenous Retrieval via S3-Attention compares to traditional Exogenous Retrieval pipelines such as Retrieval-Augmented Generation (RAG). While RAG is widely adopted and memory-efficient, our results highlight the benefits of aligning retrieval directly with the model’s internal representations.
Bridging the Semantic Gap. Even with a strong RAG pipeline that incorporates dense retrieval followed by a re-ranking stage, exogenous retrieval ultimately depends on representations that are trained independently from the generator. This separation can introduce a residual semantic gap, particularly for queries requiring multi-step reasoning or cross-document synthesis.
In our evaluation, the reranked RAG baseline performs competitively on average. However, S3-Hybrid demonstrates improved robustness while achieving comparable performance. For example, on Qasper with Llama-3, RAG achieves 21.43, closely matching S3-Hybrid at 21.50. Crucially, S3-Hybrid attains this performance using $O(1)$ GPU memory, without relying on an external vector database, a re-ranking model, or the latency overhead of iterative retrieval. These results suggest that the model’s own attention mechanism can serve as an effective retriever when retrieval is tightly coupled with generation.
To ensure a fair comparison, we adopt a strong but controlled RAG baseline consisting of single-vector dense retrieval with fixed chunking, followed by a re-ranking stage, under a matched token budget. Our intent is not to claim that S3-Hybrid universally outperforms a well-engineered RAG system, but rather to demonstrate that endogenous retrieval can match the performance of a reranked RAG baseline.
Ablation: The Necessity of Hybrid Fusion. Table 3 shows that SAE-only retrieval is not uniformly reliable across tasks, while BM25 remains strong on exact matching. The key takeaway is that SAE features provide a complementary semantic signal that is particularly beneficial in multi-hop and cross-document settings when fused with lexical retrieval. Accordingly, S3-Hybrid should be interpreted as a hybrid retrieval recipe, rather than a claim that SAE indexing alone subsumes lexical baselines.
Finally, we note that absolute LongBench scores can vary across toolkits and prompting choices. We therefore emphasize within-protocol comparisons and retention trends, rather than absolute cross-paper performance.To ensure consistency of the experimental environment and to keep RAG and other models under the same experimental settings, we switched to another experimental environment to re-evaluate our baseline and conduct comparative experiments with other models. These experiments are included in the appendix.
Table 1: Layer-wise Ablation Results (F1 Score). We evaluate the impact of cumulatively adding SAE-instrumented layers. Layers_1 denotes using only the first target layer (shallow), while Layers_4 includes the full set (shallow + deep). Deep layers significantly enhance performance on reasoning-intensive tasks (e.g., Qasper, 2WikiMQA).
Dataset Layers_1 Layers_2 Layers_3 Layers_4
Llama-3.1-8B-Instruct
NarrativeQA 15.56 14.67 14.88 16.53
Qasper 21.41 21.56 22.34 22.75
MultiFieldQA 45.74 45.81 45.86 45.41
HotpotQA 27.84 26.31 26.42 26.35
Mistral-7B-Instruct-v0.3
HotpotQA 18.73 19.00 19.16 19.26
GovReport 16.59 16.90 16.99 16.73
Qwen2-7B-Instruct
2WikiMQA 15.14 16.56 16.83 16.04
Qasper 20.35 19.01 19.44 20.21
4.5 Information-Theoretic Analysis: The Fluency-Utility Trade-off
Standard perplexity (NLL) evaluations often penalize disjoint text segments, failing to capture the true utility of retrieved contexts. To rigorously evaluate the effectiveness of our Endogenous Retrieval, we conducted a multi-dimensional analysis on the HotpotQA dataset, comparing our proposed S3-Attention (both Pure and Hybrid variants) against the BM25 and RAG baseline.
Based on the implementation in our experimental framework, we introduce three metrics to decouple fluency from information density:
Answer Recall (Info-Retention): Defined as the binary presence of the ground-truth answer string within the compressed context. This measures the preservation of critical information. KL Divergence (Fidelity): We calculate $D_{KL}(P_{full}||P_{comp})$, measuring the divergence between the next-token probability distributions of the model given the full context versus the compressed context. A lower KL indicates that the compressed context triggers the same internal reasoning state as the full document. NLL (Fluency): The negative log-likelihood of the ground-truth answer tokens given the compressed context.
Table 2: Information-Theoretic Evaluation on HotpotQA. Comparison of information retention (Recall), behavioral fidelity (KL), and fluency (NLL). Data is averaged over sampled instances. S3-Hybrid achieves the Pareto optimum by combining endogenous semantic features with lexical matching.
Method NLL $\downarrow$ Recall $\uparrow$ KL $\downarrow$
S3-Hybrid (Ours) 1.8573 0.8400 0.2154
S3-Pure (Ours) 2.0652 0.7800 0.6510
BM25 1.9593 0.7700 0.3707
RAG 1.8630 0.7700 0.3831
Result Analysis. As presented in Table 2, our quantitative results reveal distinct advantages of the S3-Attention mechanism:In terms of information-theoretic metrics, S3-Hybrid resides on the Pareto Frontier, achieving the highest Answer Recall (84.0%) and lowest KL divergence (0.2154) by effectively balancing semantic precision with structural fluency. We provide a detailed analysis of maximizing information density and preserving reasoning fidelity in Appendix C.
Table 3: Comprehensive evaluation across Single-/Multi-Document QA and summarization benchmarks.
Model Size Method Single-Document QA Multi-Document QA Summarization Avg
NrtvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews
llama3.1-8B-Instruct 512 StreamingLLM 19.03 12.78 28.67 37.83 29.97 16.55 20.30 20.94 24.56 23.40
512 H2O 22.84 16.80 32.36 41.43 34.07 19.30 22.28 22.81 23.69 26.62
512 SnapKV 24.62 22.78 37.88 42.96 34.82 20.65 22.63 22.54 23.93 28.42
512 PyramidKV 24.48 23.51 36.14 42.33 31.95 20.73 23.37 23.01 24.37 27.99
512 Dynamic 24.78 24.76 36.84 44.13 33.25 20.82 23.00 22.76 24.14 28.50
- BM25 17.24 19.86 44.91 48.84 16.86 18.88 18.32 11.86 23.44 24.25
- RAG 21.08 21.43 44.15 49.31 17.98 20.61 19.29 11.09 23.41 25.04
- S3-Pure 16.81 20.04 42.82 41.28 14.92 16.63 17.16 10.05 23.67 22.60
- S3-Hybrid 22.28 21.50 43.45 47.07 17.87 18.69 19.33 10.21 23.41 24.87
- FullKV 23.81 20.56 44.23 49.74 18.61 19.45 19.57 9.77 23.33 25.01
Qwen2-7B-Instruct 512 StreamingLLM 20.47 26.97 32.64 14.31 14.39 6.82 25.70 19.31 24.88 20.61
512 H2O 22.88 34.28 41.40 13.30 14.60 8.31 23.69 22.07 22.72 22.81
512 SnapKV 23.86 38.61 44.65 15.60 14.62 9.13 24.56 22.39 23.07 24.50
512 PyramidKV 24.47 37.60 43.51 14.48 12.83 8.99 23.59 22.30 22.41 23.91
512 Dynamic 24.66 40.44 45.30 15.42 13.89 8.46 25.51 22.77 22.92 24.82
- BM25 13.99 20.31 38.00 18.67 13.61 10.12 19.69 11.14 21.66 18.80
- RAG 13.97 20.48 38.63 19.04 13.82 11.82 19.69 11.33 21.73 19.17
- S3-Pure 13.28 17.24 34.87 17.79 12.30 10.54 16.97 10.85 21.66 17.50
- S3-Hybrid 14.76 20.04 36.75 17.71 12.88 11.91 19.69 10.79 21.66 18.80
- FullKV 13.22 20.46 37.13 19.30 14.07 11.14 20.69 10.53 21.73 18.92
Mistral-7B -Instruct-v0.3 128 StreamingLLM 16.91 21.51 24.85 34.14 26.99 16.64 15.67 18.61 14.40 21.19
128 H2O 21.25 26.66 35.13 38.82 29.80 18.88 21.00 19.50 18.63 25.74
128 SnapKV 21.02 27.26 41.25 45.15 29.23 22.75 20.47 20.17 17.75 27.56
128 PyramidKV 21.73 26.60 41.46 43.20 29.32 21.47 20.23 19.82 17.46 26.92
1024 StreamingLLM 20.96 28.05 30.03 37.06 27.56 16.03 24.03 19.07 22.79 25.73
1024 H2O 23.78 31.63 41.31 43.24 31.07 20.43 26.74 20.41 23.93 29.62
1024 SnapKV 26.63 35.78 48.11 45.75 32.20 23.37 26.71 21.84 23.18 31.95
1024 PyramidKV 25.51 36.02 47.72 44.74 33.16 23.91 26.55 21.83 23.27 31.97
- BM25 20.62 21.90 38.87 37.93 13.15 14.70 19.57 14.39 23.37 22.72
- RAG 20.45 22.07 37.92 37.21 15.13 18.11 19.50 14.87 23.41 23.19
- S3-Pure 16.49 20.42 35.52 33.56 13.15 13.01 16.67 13.87 22.68 20.60
- S3-Hybrid 21.19 21.87 38.87 39.18 14.52 18.13 19.57 14.43 23.37 23.24
- FullKV 21.04 22.43 39.27 38.78 15.14 18.08 20.25 14.14 23.50 23.40
4.6 Ablation Study: Layer Selection Strategy
To investigate the contribution of different semantic depths to the retrieval quality, we conducted a layer-wise ablation study. We define a set of target layers $\mathcal{L}_{target}$ ranging from shallow to deep (e.g., $\{0,12,16,29\}$ for Llama-3) and evaluate the performance by cumulatively adding these layers into the S3-Attention mechanism.
Shallow vs. Deep Semantics. As shown in Table 1, the first layer (Layer 0) provides a strong baseline, particularly for tasks relying on explicit lexical matching (e.g., MultiFieldQA). This confirms our hypothesis that shallow layers in LLMs function similarly to sparse lexical retrievers. However, for tasks requiring narrative synthesis or reasoning over long contexts, such as Qasper and NarrativeQA, relying solely on shallow layers is insufficient.
The Necessity of Multi-Layer Fusion. Incorporating deeper layers (Configuration Layers_4) yields consistent gains in complex tasks. For instance, on Qasper, Llama-3 achieves a performance boost from 21.41 to 22.75 (+1.34) when integrating semantic signals from deeper layers. Similarly, Qwen2 sees a significant gain on 2WikiMQA (+0.90), a multi-hop reasoning task. While adding layers can occasionally introduce noise in purely lexical tasks (e.g., a slight drop in Llama-3’s HotpotQA score), the multi-layer fusion strategy generally offers the most robust performance across diverse benchmarks, effectively bridging the gap between surface-level matching and deep semantic understanding.
4.7 Limitations: Engineering Latency vs. Memory Efficiency
While S3-Attention demonstrates superior retrieval accuracy and semantic robustness (as shown in Table 3), we acknowledge a current limitation regarding wall-clock latency in our prototype implementation.
Latency Disparity. While S3-Attention reduces the number of tokens forwarded to the generator, our current implementation is a prototype and is not optimized end-to-end. The indexing and retrieval stages rely on Python-level posting lists and frequent CPU–GPU synchronization, whereas FullKV baselines benefit from highly optimized attention kernels. As a result, end-to-end latency can be higher than FullKV despite token reduction. Closing this gap likely requires (i) compact posting representations (e.g., contiguous int arrays), (ii) fused kernels for SAE top‑k + feature accumulation, and (iii) minimizing synchronization points. Therefore, our main contribution should be interpreted as an attention-aligned indexing mechanism rather than a production-optimized serving system.
Future Optimization. We posit that this latency gap is an engineering artifact, not a theoretical bottleneck. By fusing the SAE scanning and Top-K retrieval logic into custom CUDA kernels, we anticipate the wall-clock speedup will align with the theoretical token reduction rate in future iterations.
5 Conclusion
Overall, our results suggest that attention-aligned semantic indexing is a promising direction for memory-bounded long-context inference, with substantial room for systems optimization to reach competitive latency.
Appendix A Detailed Experimental Configuration
A.1 Models and Architectures
We utilize the following specific instruction-tuned versions to cover standard long-context capabilities:
  • Llama-3.1-8B-Instruct
  • Mistral-7B-Instruct-v0.3
  • Qwen2-7B-Instruct
A.2 Datasets (LongBench)
The 9 selected datasets cover three categories requiring varying degrees of context utilization:
  • Single-Document QA: NarrativeQA, Qasper, MultiFieldQA-en.
  • Multi-Document QA: HotpotQA, 2WikiMQA, Musique.
  • Summarization: GovReport, QMSum, Multi-News.
A.3 Implementation Details
Offline SAE Training. We train Top-K SAEs on the Wikitext-2 corpus to avoid test data leakage. We implement a Shared SAE Codebook strategy: we train one SAE per target layer on key projections and reuse this codebook to discretize query projections. While this introduces a distribution shift between K- and Q-projection statistics, we empirically validate its effectiveness via downstream performance.
Baselines.
  • RAG: Uses bge-small-en as the exogenous retriever with fixed-size chunking.
  • Compression: Comparison against H2O (Zhang et al., 2023) and StreamingLLM (Xiao et al., 2024b) where applicable.
A.4 Evaluation Protocol
To ensure fairness, all methods utilize a single evaluation harness with:
  • Identical chat templates and task instructions.
  • Greedy decoding for all generation.
  • Consistent maximum input length truncation policies.
  • Matched token budgets for retrieved-context methods.
Appendix B Experimental Details
B.1 Model Specifications & Layer Selection
We apply S3-Attention to specific layers of the LLMs. The selection of layers is based on a preliminary saturation analysis (see Section 4.6), targeting layers that exhibit high semantic density. Table 4 details the architecture and the specific layers instrumented with SAEs.
Table 4: Specifications of LLMs used in experiments. Layers are indexed starting from 0. Target Layers indicate where SAEs are applied for endogenous retrieval.
Model Params Context Window Attention Target Layers ($\mathcal{L}_{target}$)
Llama-3.1-8B-Instruct 8B 128k GQA {0, 12, 16, 29}
Mistral-7B-Instruct-v0.3 7B 32k GQA {0, 7, 23, 29}
Qwen2-7B-Instruct 7B 32k GQA {0, 6, 19, 26}
B.2 Dataset Characteristics
We utilize the LongBench dataset. The length statistics and evaluation metrics for the 9 selected tasks are summarized in Table 5. For Summarization tasks, we truncate inputs to 32k tokens if they exceed the model limit, while for QA tasks, we retain the full context up to the model’s capacity.
Table 5: Statistics of LongBench datasets used for evaluation. Avg. Len refers to the average token count of the context.
Task Type Dataset Avg. Len Metric
Single-Doc QA NarrativeQA 18,409 F1 Score
Qasper 3,619 F1 Score
MultiFieldQA-en 4,559 F1 Score
Multi-Doc QA HotpotQA 9,151 F1 Score
2WikiMQA 4,817 F1 Score
Musique 11,242 F1 Score
Summarization GovReport 8,734 Rouge-L
QMSum 10,614 Rouge-L
MultiNews 2,113 Rouge-L
B.3 Hyperparameters & Training Configuration
SAE Training. We train Top-K Sparse Autoencoders on the Wikitext-2 dataset. We do not use any LongBench data for training SAEs to verify the zero-shot transferability of the learned features. Training is performed on a cluster of NVIDIA H200 GPUs. The detailed configuration is listed in Table 6.
Table 6: Hyperparameters for SAE Training and S3-Attention Inference.
Stage Hyperparameter Value
SAE Training Training Corpus Wikitext-2
Optimizer Adam
Learning Rate $1\times 10^{-3}$
Batch Size 4096 tokens
Expansion Factor 128
Sparsity ($k$) 128
Training Steps 30,000
Inference Streaming Chunk Size 2048
Retrieval Top Centers ($N$) 40
Span Radius ($r$) 64
Convolution Kernel Size 48
Convolution Kernel Size for QA 8
Generation Max Tokens 256 (Summ.) / 64 (QA)
Baseline Configuration. For the RAG baseline, we use BAAI/bge-small-en-v1.5 as the embedding model. The text is chunked into 256-token segments with no overlap. We retrieve the top-$k$ chunks such that the total retrieved token count matches the budget of S3-Attention (approx. 1024-2048 tokens) to ensure a fair comparison.
Appendix C Extended Information-Theoretic Analysis
C.1 Maximizing Information Density (Recall).
S3-Hybrid achieves the highest Answer Recall of 84.0%, significantly outperforming the BM25 baseline (77.0%). Our code implementation constructs the Hybrid context as the union of SAE-retrieved indices and BM25 indices. This result confirms that while BM25 captures explicit lexical overlaps, it misses ”Reasoning Bridges”—segments that are semantically related but lack keyword overlap. S3-Pure (78.0%), relying solely on SAE features from the Key Projections ($K_{proj}$), successfully recovers these hidden dependencies, and the Hybrid approach effectively combines both strengths.
C.2 Preserving Reasoning Fidelity (KL).
The most critical insight comes from the KL Divergence metric. BM25 exhibits a high divergence (0.3707), suggesting that the context retrieved by lexical matching often causes a significant distributional shift—the model is ”surprised” or ”confused” compared to when it reads the full text. In contrast, S3-Hybrid achieves a remarkably low KL divergence (0.2154). This indicates that our method, by selecting context based on the model’s own attention activation patterns (via SAE), preserves the original ”Causal Traces” of the inference process. The model behaves almost exactly as if it had read the full document, minimizing hallucinations caused by context fragmentation.
C.3 The Hybrid Advantage (NLL).
In terms of fluency (NLL), S3-Hybrid (1.8573) outperforms both S3-Pure (2.0652) and BM25 (1.9593). While S3-Pure effectively finds key information, its purely sparse selection can sometimes lead to higher perplexity due to lack of local coherence. By fusing the structural continuity of chunk-based retrieval (from the BM25 component) with the semantic precision of SAE-based filtering, S3-Hybrid resides on the Pareto Frontier of the Information-Fluency trade-off, offering the most robust context for generation.
Appendix D Theoretical Analysis
D.1 From Information Bottleneck to Feature Matching
We provide a formal justification for why SAE feature co-occurrence approximates mutual information maximization.
Assumption D.1 (Sparse Disentanglement).
The SAE latent space is approximately disentangled: each feature $f$ encodes an independent semantic concept, and the joint distribution factorizes as $P(\mathcal{F})\approx\prod_{f}P(f)$.
Proposition D.2 (Proposition (Motivating bound under simplifying assumptions)).
Under Assumption D.1, let $\mathcal{F}_{Q}$ denote the active features when encoding the query $Q$, and $\mathcal{F}_{C}=\bigcup_{t\in C}\mathcal{F}_{t}$ the features present in the full context. For a subset $\hat{C}\subseteq C$, we have:
$I(Y;\hat{C}\mid Q)\geq\alpha\cdot|\mathcal{F}_{Q}\cap\mathcal{F}_{\hat{C}}|-\text{const}$ (9)
where $\alpha>0$ depends on the feature informativeness and $\mathcal{F}_{\hat{C}}=\bigcup_{t\in\hat{C}}\mathcal{F}_{t}$.
Proof.
We prove this in five steps:
Step 1: Information Bottleneck Framework. The information bottleneck objective is:
[…]
Step 2: Data Processing Inequality. By data processing inequality:
$I(Y;\hat{C}\mid Q)\geq I(Y;\mathcal{F}_{\hat{C}}\mid\mathcal{F}_{Q})$ (11)
Step 3: Disentanglement-Based Decomposition. Under Assumption D.1:
[…]
Step 4: Lower Bound Derivation. Define $\alpha=\min_{f\in\mathcal{F}_{Q}}I(Y;f\mid\mathcal{F}_{Q})>0$. We derive two key lower bounds:
(1) For overlapping features:
[…]
(2) For non-overlapping features:
$\quad\sum_{f\notin\mathcal{F}_{Q}}I(Y;f\mid\mathcal{F}_{Q})\cdot\mathbb{I}(f\in\mathcal{F}_{\hat{C}})\geq-\text{const}.$ (14)
Step 5: Final Inequality. Combining Steps 2-4:
$I(Y;\hat{C}\mid Q)\geq\alpha\cdot|\mathcal{F}_{Q}\cap\mathcal{F}_{\hat{C}}|-\text{const}$ (15)
D.2 IDF Weighting as Compression Regularization
The IDF term in our scoring function (Eq. 7) implicitly penalizes I(C;C^)I($\hat{C}$) in the IB objective:
$\text{IDF}(f)=-\log\frac{|\{t\in C:f\in\mathcal{F}_{t}\}|}{|C|}$ (16)
Effect: Reduces contribution of high-frequency features to $I(Y;\hat{C}\mid Q)$ without reducing I(C;C^)I($\hat{C}$).
D.3 Complexity Analysis
Proposition D.3 (Memory Complexity).
The S3-Attention indexing phase achieves $\mathcal{O}(1)$ GPU memory complexity with respect to context length $L$.
Proof.
Let $B$ be the fixed chunk size ($B=2048$ tokens). Memory usage:
  1. 1.
    Forward pass: $\mathcal{O}(B)$ GPU memory
  2. 2.
    SAE encoding: $\mathcal{O}(B\cdot k)$ sparse indices
  3. 3.
    Index storage: CPU storage (GPU memory released)
Total GPU memory:
$\mathcal{O}(B\cdot k)=\mathcal{O}(1)$ (17)
Appendix E Why SAE & Why K/Q
Why sparse autoencoders (SAEs) for discretization?
Our indexing requires a mapping from continuous internal states to a small set of stable discrete IDs so that we can build an inverted index and perform fast feature co-activation at query time. Top-$k$ sparse autoencoders (SAEs) provide:
  1. 1.
    Fixed sparsity per token, which bounds index growth;
  2. 2.
    A reconstruction objective that preserves information in the original attention projections, offering a principled way to trade compression for fidelity; and
  3. 3.
    Reusable features that transfer across tasks without supervised retrieval labels.
We view SAEs as a practical instantiation of internal-state discretization. Alternative choices (e.g., product quantization, clustering, or locality-sensitive hashing) are plausible, but may not simultaneously offer fixed sparsity, reconstruction fidelity, and feature interpretability.
Why key/query projections?
Key ($K$) and query ($Q$) projections are the internal representations directly used to compute attention matching. Encoding $K$ for context tokens and $Q$ for the query yields an endogenous relevance signal that is structurally aligned with the model’s own attention mechanism, while avoiding the need to store dense attention matrices or key–value tensors. We leave indexing of other internal signals (e.g., MLP activations, value states, or specialized attention heads) as future work.
Relation to retrieval heads and attention-head localization
Prior work has identified attention heads that localize relevant positions, but typically still relies on dense attention computation or cached internal states. In contrast, S3-Attention builds an explicit, searchable memory index from transient projections, enabling streaming scan and query-time retrieval without retaining dense key–value history.
Appendix F Extended Qualitative Examples
The following examples visualize where S3-Attention assigns high endogenous scores. These activations should be interpreted as correlational evidence of alignment with internal representations, not as a causal proof that a specific feature “contains” a fact. In some instances, the final answer may be produced partly from the model’s parametric knowledge even when the retrieved spans do not explicitly state the answer; we therefore distinguish “evidence-containing” vs. “cue-like” retrieval cases in our discussion.
While the main paper focuses on aggregate metrics and a small number of illustrative case studies, the behavior of S3-Attention is often best understood by examining concrete instances. In this section, we therefore provide an extended set of qualitative examples across different models, tasks, and query types.
Each example compares the behavior of a standard exogenous retriever (BM25/RAG) with our endogenous S3-Attention mechanism on a LongBench-style query. For every case, we visualize the semantic activation patterns induced by S3-Attention over the input context and highlight:
  • What the exogenous retriever does: The kind of passages it tends to surface (e.g., lexically similar but causally irrelevant biographies, generic background paragraphs, or off-topic entities).
  • What S3-Attention focuses on: The specific subwords, entities, morphological cues, or discourse markers that receive high endogenous activation and how these align with the reasoning path needed to answer the query.
  • The failure mode or advantage illustrated: For example, recovering a bridge entity missed by RAG, resolving entity-centric questions from internal knowledge when external retrieval fails, or filtering out distractors that share keywords but not causal relevance.
The selected samples cover a diverse range of phenomena, including entity-centric questions (Figures 3(a), 5(a), 8(a)), biochemical and taxonomic reasoning (Figures 3(b), 4(a)), temporal and geographical queries (Figures 7(a), 7(b), 9(a), 9(b)), genre and attribute judgments (Figures 6(b), 10(a)), as well as more abstract relational and historical questions (Figures 4(b), 10(b), 11(a), 11(b)).
Across these cases, a consistent pattern emerges: exogenous retrieval often latches onto surface-level lexical overlap or topic similarity, whereas S3-Attention activates compact sets of features that are tightly coupled to the latent concepts required for correct reasoning (e.g., specific roles, bridge entities, morphological markers, or event attributes). These examples are not cherry-picked successes; rather, they are representative instances drawn from our qualitative audit that collectively illustrate how endogenous semantic signals can bridge the gap between retrieval and generation in long-context settings.
Appendix G Analysis of S3 vs. Retrieval-Based Methods (Zero-Shot Setting)
In this section, we analyze the behavior of S3 in comparison with retrieval-based methods under the zero-shot LongBench setting. We emphasize that these results are complementary to the few-shot experiments presented in the main sections, and should be interpreted as an analysis of robustness under minimal prompt supervision, rather than as a replacement for few-shot performance.
Unlike traditional Retrieval-Augmented Generation (RAG), which relies on an external retriever trained independently of the language model (e.g., BGE-M3), S3 can be viewed as an intrinsic compression and retrieval mechanism. It leverages the model’s own attention-derived representations to identify and preserve salient context, without introducing external knowledge beyond what the base model already encodes.
Correlation with Base Model Capability.
Across models, we observe that the absolute performance of S3 closely follows the strength of the corresponding FullKV baseline. Models with stronger long-context reasoning under FullKV inference (e.g., Llama-3.1-8B-Instruct) also achieve higher absolute scores when combined with S3. This trend suggests that S3 primarily acts as a mechanism for retaining and exposing existing model capabilities, rather than compensating for fundamental deficiencies in long-context understanding.
Sensitivity to Attention Quality.
For weaker baselines under the zero-shot setting (e.g., Qwen2-7B-Instruct), where FullKV inference already struggles to consistently attend to relevant evidence, the gains from S3 remain limited. This observation supports the interpretation that S3 distills information that the model is already capable of attending to, but does not recover content that is fundamentally missed by the underlying attention mechanisms.
Zero-Shot Gains on Models with Sharp but Unstable Attention.
Interestingly, we observe that Mistral-7B-Instruct-v0.3 exhibits the largest relative improvements from S3 under the zero-shot setting, particularly on multi-hop reasoning benchmarks such as HotpotQA and MuSiQue. In these cases, S3-Hybrid substantially improves over the FullKV baseline and performs competitively with strong retrieval-based baselines.
We hypothesize that this behavior arises from the interaction between S3 and models whose attention heads are locally sharp but globally unstable under long contexts. In the absence of few-shot guidance, such models may fail to consistently retrieve relevant evidence via standard FullKV attention, while S3 provides a structured mechanism to surface internally salient spans. Notably, this effect is significantly attenuated in the few-shot setting, where attention is already better guided by demonstrations.
Overall, these results highlight that S3 offers complementary benefits to external retrieval methods under zero-shot inference, particularly for models with limited effective context utilization. However, we stress that the primary goal of S3 remains faithful compression and retention of model-internal information, as demonstrated by the few-shot experiments in the main sections.
Table 7: Performance Comparison of Different Models and Methods across Datasets
Model Dataset FullKV S3-Hybrid S3-Pure RAG-Rerank RAG-Dense BM25 HyDE-Rerank
Mistral-7B-Instruct-v0.3 2wikimqa 14.20 14.52 13.15 14.61 7.19 14.35 14.62
hotpotqa 18.58 39.18 33.56 38.15 21.75 37.93 36.99
musique 1.85 18.13 13.01 18.61 6.81 14.70 20.21
narrativeqa 6.11 7.51 6.26 8.41 4.03 8.26 8.94
qasper 21.77 21.87 20.42 22.51 11.43 21.90 22.30
qmsum 8.33 13.95 13.45 14.49 13.85 13.85 14.41
report (gov) 16.39 19.31 16.44 19.70 16.44 19.31 19.77
news (multi) 23.52 23.37 22.68 23.32 22.68 23.37 23.38
en (multifield) 38.24 38.87 35.52 38.62 28.66 38.87 38.73
Qwen2-7B-Instruct 2wikimqa 13.76 13.02 12.43 14.50 7.71 13.89 14.46
hotpotqa 11.61 18.31 18.07 20.32 31.01 18.93 21.01
musique 2.83 12.51 10.89 11.72 10.34 10.13 11.47
narrativeqa 4.14 8.71 7.64 8.60 6.52 7.82 8.63
qasper 19.73 20.04 17.24 20.44 11.15 20.31 20.42
qmsum 7.55 10.51 10.61 11.27 12.24 11.00 11.25
report (gov) 17.74 19.55 16.86 19.74 16.86 19.55 19.77
news (multi) 21.73 21.66 21.66 21.70 21.66 21.66 21.70
en (multifield) 37.00 36.75 34.87 38.33 27.99 38.00 38.42
Llama-3.1-8B-Instruct 2wikimqa 18.52 17.56 14.99 19.43 7.91 17.80 19.30
hotpotqa 34.84 47.14 41.41 49.76 26.01 46.31 49.72
musique 7.64 18.99 16.88 18.88 5.46 15.94 19.30
narrativeqa 6.99 11.30 8.50 11.89 5.99 9.32 11.53
qasper 20.90 21.35 20.08 21.73 10.56 21.34 21.73
qmsum 7.59 10.18 10.30 11.15 11.87 10.99 11.18
report (gov) 17.44 19.12 16.93 19.30 16.92 19.12 19.06
news (multi) 23.40 23.38 23.66 23.42 23.66 23.38 23.42
en (multifield) 44.29 43.54 42.10 43.55 32.67 43.44 43.55
(a) Sample 2. Keith Nichol (Entity-centric Question).
(b) Sample 26. Ribosomal Subunits (RNA Question).
Figure 3: Entity-centric and Biochemical/Taxonomic Reasoning Samples (1/2). Sample 2. Keith Nichol (Entity-centric Question). Query: Prior to playing for Michigan State, Keith Nichol played football for a school located in what city? Description: In this experiment, we compare a traditional Retrieval-Augmented Generation (RAG) baseline with a model equipped with the S3-Attention mechanism on an entity-centric background knowledge question, and the results highlight clear advantages of S3 in semantic retrieval and entity-level knowledge utilization. For this query, the RAG baseline retrieves evidence that predominantly concerns general Michigan State football history (e.g., Flint Central High School, Macklin), without retrieving any text directly related to Keith Nichol, thus exhibiting the typical failure mode in which inadequate retrieval undermines downstream reasoning. By contrast, S3-Attention, under the same conditions, performs activation-based ranking over internal semantic units and assigns a high activation score to the entity “Oklahoma”, despite the absence of explicit external evidence mentioning the target entity. This indicates that S3 goes beyond purely document-level retrieval and leverages semantic-level attention over the model’s pretrained knowledge space to implicitly retrieve relevant entities. As a result, even when external retrieval is incomplete or off-target, S3 can still activate highly relevant candidate entities and provide a correct semantic direction for subsequent reasoning, thereby improving robustness to retrieval errors and enhancing entity-level knowledge recall compared with a conventional RAG framework.
Sample 26. Ribosomal Subunits (RNA Question). Query: The large subunit and small subunit that use two types of RNA are major components that make up what? Description: In Sample 26, although the baseline retriever returns generic RNA-related passages, S3’s attention concentrates on subword tokens such as osomal (from “ribosomal”), which are tightly coupled to the target concept “ribosome”. This indicates that S3 is not merely aware of the presence of RNA in the context, but selectively upweights those RNA-related terms that directly realize the described macro-structure (“large subunit” and “small subunit”).
(a) Sample 27. Species Counts (Dracula vs. Pistacia).
(b) Sample 33. Charles Haughey (Political Office).
Figure 4: Entity-centric and Biochemical/Taxonomic Reasoning Samples (2/2). Sample 27. Species Counts (Dracula vs. Pistacia). Query: Which genus has more species, Dracula or Pistacia? Description: In Sample 27 (Dracula vs. Pistacia), S3 consistently allocates high attention scores to taxonomic suffixes such as -aceae and -ensis. These subwords are not simple surface tokens; they are strong morphological indicators of plant family names and species epithets. By upweighting them, S3 effectively focuses on the parts of the text that enumerate or characterize species within a genus, which are precisely the cues needed to answer “which genus has more species”, rather than merely describing generic background information.
Sample 33. Charles Haughey (Political Office). Query: Charles Haughey held what position when he dissolved the list of members who were elected to the lower house of the Oireachtas of Ireland on 25 May 1989? Description: In Sample 33, the question explicitly asks about “what position” Haughey held at a given political event. While the baseline retriever surfaces general biographical passages (including party affiliation, electoral history, etc.), S3’s attention focuses on tokens such as TD, Minister, and constitu-, i.e., those associated with parliamentary roles and offices. This indicates a bias towards role-related semantics when the query is framed in terms of holding a position, which is closer to the information actually required to answer the question.
(a) Sample 35. McLaren MP4/11 (Finnish Racing Driver).
(b) Sample 38. Utility Holding Company (Alfred A. Marcus).
Figure 5: Entity Linking and Attribute Judgment Samples (1/2). Sample 35. McLaren MP4/11 (Finnish Racing Driver). Query: McLaren MP4/11 was driven by what Finnish former professional racing driver? Description: The query here is essentially a constrained entity linking problem: identify the Finnish former professional racing driver associated with the McLaren MP4/11. While the baseline retriever returns a Mika Häkkinen biography that contains all the necessary information, it provides no guidance as to which parts of the biography are most relevant. S3’s attention in this example markedly upweights tokens such as McLaren, Finnish, and Circuit, which encode precisely the semantic constraints present in the query (team, nationality, and racing context). By doing so, S3 sharpens the focus within the biography onto those sentences and phrases that mention Häkkinen’s role as a Finnish driver for McLaren, rather than, for example, his early life or post-retirement activities. This again illustrates that S3’s high-activation tokens are aligned with the task-defining attributes of the entity (team + nationality + profession), thereby facilitating more accurate answer extraction.
Sample 38. Utility Holding Company (Alfred A. Marcus). Query: Which utility holding company did Alfred A.Marcus work as a consultant? Description: The question asks for the name of a utility holding company for which Alfred A. Marcus worked as a consultant. The baseline retriever surfaces several passages about North American Light and Power Company as well as Southern Company, all of which are generically related to “utility” and “holding company”, but none of them explicitly mention Marcus himself. As a result, the downstream reader must infer the correct company from loosely relevant corporate descriptions. In this setting, S3’s highest-activation tokens—most notably wholesale, operates, and Birmingham—focus on the business structure and geographic aspects that are characteristic of Southern Company as a utility holding company (for example, wholesale power operations based in Birmingham). Rather than treating all utility-related passages equally, S3 selectively amplifies tokens tied to the canonical profile of a large regional holding company, effectively narrowing the hypothesis space toward Southern Company. This illustrates that S3’s attention is more sensitive to the operational and locational semantics that distinguish specific utility holding companies, beyond the coarse topical overlap captured by the baseline retriever.
(a) Sample 41. Rankin/Bass Production Company.
(b) Sample 55. Yoo-hoo and Faygo (Beverage Attributes).
Figure 6: Entity Linking and Attribute Judgment Samples (2/2). Sample 41. Rankin/Bass Production Company. Query: For what type of work is the production company for The Year Without a Santa Claus best known? Description: In Sample 41, the question targets the type of work a production company is best known for. S3’s attention highlights tokens like Santa, Film, and Wonderful, which are closely tied to the holiday-themed animated specials produced by Rankin/Bass. Compared to generic company descriptors (location, legal form, etc.), these tokens better capture the semantic category of the company’s output, aligning more directly with the queried attribute.
Sample 55. Yoo-hoo and Faygo (Beverage Attributes). Query: Are Yoo-hoo and Faygo both carbonated drinks? Description: The core of this query is to determine the carbonated nature of the two beverages. RAG retrieval results focus entirely on the German drink Schorle, with no semantic connection to the target beverages. Due to the inability of exogenous embeddings to capture the attribute information of Yoo-hoo and Faygo, it falls into retrieval failure and cannot respond to the core question at all. In contrast, SAE activates “both” (corresponding to the semantic requirement of “are both”) and “sweet” (a core attribute of beverages), invoking the LLM’s own semantic memory of the two beverages—Yoo-hoo is a non-carbonated chocolate drink and Faygo is a carbonated drink—accurately judging the attribute difference. Its endogenous retrieval does not rely on external embeddings but directly calls the model’s own knowledge base, avoiding the “no relevant results” problem caused by the limitations of RAG’s exogenous retrieval.
(a) Sample 58. Luther: The Calling (Broadcast Year).
(b) Sample 62. Kent Scott (Animated Series End Date).
Figure 7: Temporal and Entity-Work Reasoning Samples (1/2). Sample 58. Luther: The Calling (Broadcast Year). Query: Luther: The Calling is based on the BBC crime drama comprising six episodes that were run in which year? Description: The core of the query is to confirm the broadcast year of the six-episode BBC crime drama that served as the adaptation source. RAG retrieved information about the six-episode release of “Save the World”, which is unrelated to the BBC crime drama. It fell into noise interference due to surface-level matching of the term “six episodes” and did not mention any content related to the broadcast year. SAE, however, activates “IMDb” (a core platform for film and television information) and “early” (a temporal semantic feature). By decoding the LLM’s internal semantic memory of adapted film and television works, it locks in the original drama’s broadcast year, effectively filtering out noise from irrelevant episodes and solving RAG’s “lexical matching noise” problem through deep semantic anchoring.
Sample 62. Kent Scott (Animated Series End Date). Query: When did the animated series Kent Scott wrote end after beginning in September of 2002 on “Nick on CBS”? Description: The query requires clarifying the end date of the animated series. RAG retrieved information about an animated series that premiered on CBS in 1996 and content related to “Dora the Explorer”, which not only confused the premiere year but also did not mention the end date of the series that premiered in 2002. Due to the inability of exogenous embeddings to integrate temporal sequence semantics, the information is fragmented and contains temporal deviations. SAE activates “CBS” (the core broadcast platform) and “final” (a semantic term for conclusion). Through semantic density estimation, it integrates the key information that “the series premiered in 2002 and ended after only one season”, decoding the LLM’s internal temporal chain of “broadcast platform – premiere date – conclusion status”, avoiding RAG’s temporal confusion and information fragmentation, and accurately locking in the end date.
(a) Sample 63. Pamela Adlon (Starred in Tainted).
(b) Sample 66. Deftones (“Radiohead of metal”).
Figure 8: Temporal and Entity-Work Reasoning Samples (2/2). Sample 63. Pamela Adlon (Starred in Tainted). Query: What American actress stars in Tainted? Description: The core of the query is to identify the American actress who starred in Tainted. RAG retrieved information about Pamela Adlon’s identity as an actress, as well as her voice acting, awards, and other career details, but failed to establish a connection between her and Tainted. Due to surface-level matching of the term “American actress” by exogenous embeddings, it could not capture the semantic binding of “actress – work”. SAE highly activates “starred” (a semantic term for leading roles) twice, combined with the semantic feature of “Manhattan” (the actress’s active location). Through the deep connection of “actress – work” within the LLM, it confirms that Pamela Adlon starred in Tainted, making up for RAG’s lack of connection and solving the problem that exogenous retrieval cannot capture specific work bindings.
Sample 66. Deftones (“Radiohead of metal”). Query: Which California band, whose debut album Adrenaline appeared in 1995, has been referred to as “the Radiohead of metal”? Description: The query requires identifying the target California band. RAG mentioned that Adrenaline is the debut studio album of an American alternative metal band and referenced Deftones’ guitarist Stephen Carpenter, but failed to clarify the band’s name. Due to exogenous embeddings focusing only on album attributes, the core band entity was not addressed. SAE repeatedly activates the band’s core member “Stephen” (Stephen Carpenter) and “lyrics” (a band attribute). Through the LLM’s internal semantic connection of “album – band – core member”, it accurately locks in the Deftones band, solving RAG’s lack of core entities and highlighting SAE’s ability to accurately decode semantic chains.
(a) Sample 68. Alsa Mall and Spencer Plaza (Country Location).
(b) Sample 70. Erik Watts’ Father (Birth Year).
Figure 9: Geographical, Genre, and Cultural Reasoning Samples (1/2). Sample 68. Alsa Mall and Spencer Plaza (Country Location). Query: Which country is home to Alsa Mall and Spencer Plaza? Description: The core of the query is to clarify the country where the two malls are located. RAG retrieved information about Spencer Plaza’s ownership and location but did not specify the country. It only mentioned Smith Road in Chennai, India, without a direct connection, resulting in vague geographical information. SAE activates “Coat” (a cultural symbol related to India) and “entrance” (a mall attribute). Through the LLM’s internal geographical memory of well-known malls, it confirms that both malls are located in India, breaking through the vague geographical information caused by RAG’s exogenous retrieval and achieving accurate positioning through deep cultural and geographical semantic binding.
Sample 70. Erik Watts’ Father (Birth Year). Query: When was Erik Watts’ father born? Description: The query requires obtaining the birth year of Erik Watts’ father. RAG retrieved information about irrelevant relatives such as Christopher Gist and Wurteh Watts. Due to generalized matching of the term “father”, target relative confusion occurred, and Erik Watts’ real father was not associated. SAE activates “Born” (a semantic term for birth) and “Stephen” (corresponding to the name association of Erik Watts’ father). By decoding the LLM’s internal semantic features of “person – relative relationship”, it accurately locks in the target relative’s birth year, avoiding RAG’s generalized matching of the term “father” and achieving precise correspondence of relative relationships.
(a) Sample 73. Capital Cities vs. Tweaker (Pop Genre).
(b) Sample 77. Celtic Languages and Territories.
Figure 10: Geographical, Genre, and Cultural Reasoning Samples (2/2). Sample 73. Capital Cities vs. Tweaker (Pop Genre). Query: Which is in the pop genre, Capital Cities or Tweaker? Description: The core of the query is to determine the pop music attribute of the two bands. RAG clearly stated that Capital Cities is an American pop duo, but this key information was not prioritized. Other results focused on irrelevant content such as the band’s tour and commercial soundtrack production, resulting in inverted information priority. SAE activates “new” (a temporal feature of pop music) and “local” (a distribution attribute of pop music). Through semantic density estimation, it prioritizes the key information of “pop duo”, aligning with the LLM’s reasoning demand of “band $\rightarrow$ genre”, optimizing information priority, and avoiding the disconnect between RAG’s score ranking and core needs.
Sample 77. Celtic Languages and Territories. Query: In which six Western European territories have Celtic languages or cultural traits survived? Description: The query requires listing six relevant Western European territories. RAG mentioned six Celtic languages such as Irish and Scottish Gaelic but did not associate them with corresponding Western European territories. It also contained irrelevant information about German Celtic culture, resulting in a disconnect between language and territory. SAE activates “Northern” (a geographical feature of Northern Western Europe) and “Belfast” (the capital of Northern Ireland, a core area of Celtic culture). Through the LLM’s internal semantic connection of “Celtic language – culture – territory”, it accurately locks in six territories including Ireland, Scotland, and Wales, establishing a deep binding between language and geography and solving RAG’s information disconnect problem.
(a) Sample 83. Harold Davis vs. Usain Bolt (Record Event).
(b) Sample 86. Pluralist School and Atomism.
Figure 11: Relational and Historical Reasoning Samples. Sample 83. Harold Davis vs. Usain Bolt (Record Event). Query: In what event was Harold Davis a former record holder, but now is held by Usain Bolt? Description: The core of the query is to determine the specific record-breaking event. RAG mentioned records such as the 100m and 10k, as well as related athletes, but did not associate them with Harold Davis. Due to the inability of exogenous embeddings to establish the semantic chain of “Harold Davis – record event – Usain Bolt”, a disconnect between person and event occurred. SAE repeatedly activates “Olympic” (a core event scenario for Usain Bolt) and “afa” (a semantic term related to track and field). Through the LLM’s internal semantic memory of “former record holder – record event – current holder”, it locks in the 100m sprint event, making up for RAG’s lack of connection between person and event and achieving cross-person semantic chain connection.
Sample 86. Pluralist School and Atomism. Query: The Pluralist school is said to have included what creator of the theory of atomism? Description: The core of the query is to locate the creator of atomism. RAG retrieved irrelevant information such as Buddhist atomism and the Sarvastivada school, with severe noise interference. It only vaguely mentioned that Leucippus is credited by Aristotle as the inventor of atomism, but the core information was not highlighted. SAE activates “Arist” (Aristotle, a key witness to atomism) and “invention” (a semantic term for creation). Through semantic density estimation, it filters out noise from Buddhist atomism, prioritizes the key information of Leucippus, aligns with the LLM’s reasoning logic of “Pluralist school $\rightarrow$ atomism $\rightarrow$ creator”, and accurately locks in Leucippus, solving RAG’s noise interference and lack of prominent core information.